1. Jane's brother Andrew leaves home for school at 8:00 am. He walks at 4.2 kph. At 8:20 am Jane discovers that Andrew has left his homework at home. She decides to follow him on her bicycle and give him his homework, but she wants to be back home in time for an online test at 9:00 am. What is the minimum constant speed at which she needs to ride?

2. A car accelerates with a constant acceleration of 2.1 m/ s^2 along a long straight road. How long does it take the car to accelerate from 100 kph to 110 kph, and how far does it travel while doing so? [ take the speeds as three significant figures.]

3. A (field) hockey player is running north at a speed of 8 m/s . A ball is coming directly towards her from the north at a speed of 18 m/s. She swings her stick forward, directly north, at a speed of 6 m/s measured in her own frame of reference. What is the speed of the ball with respect to the hockey stick?
1
Expert's answer
2020-02-27T10:11:42-0500

1.


"d=vt=4.2\left(t+\frac{20}{60}\right)"

We have:


"2t=1\ h"

Thus,


"0.5v=4.2\left(0.5+\frac{20}{60}\right)"

"v=7\frac{km}{h}"

2.


"t=\frac{v_f-v_i}{a}=\frac{110-100}{3.6(2.1)}=1.3\ s"

3.


"v_y=8+18=26\frac{m}{s}"

"v=\sqrt{v_x^2+v_y^2}=\sqrt{6^2+26^2}=26.7\frac{m}{s}"